# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

### v1

@@ 6 Conclusion
This paper proves that the randomized communication complexity of permutation-invariant Boolean functions is at most quadratic of the quantum communication complexity (up to a logarithmic factor). Our results suggest that symmetries prevent exponential quantum speedups in communication complexity, extending the analogous research on query complexity. Furthermore, we prove that the Log-rank Conjecture and Log-approximate-rank Conjecture hold for non-trivial permutation-invariant Boolean functions (up to a logarithmic factor). There are some interesting problems to explore in the future.
  • Permutation invariance over higher alphabets. In this paper, the permutation-invariant function is a binary function. The interesting question is to generalize our results to larger alphabets, i.e., to permutation-invariant functions of the form $f:{X}^{n}\times{Y}^{n}\to R$ where ${X},{Y}$ and $R$ are not necessarily binary sets.
  • Generalized permutation invariance. It is possible to generalize our results for a larger class of symmetric functions. One candidate might be a class of functions that have graph-symmetric properties. Suppose $\mathcal{G}_{A},\mathcal{G}_{B}$ are two sets of $n$-vertices graphs, and $G_{n}$ is a group that acts on the edges of an $n$-vertices graph and permutes them in a way that corresponds to relabeling the vertices of the underlying graph. A function $f:\mathcal{G}_{A}\times\mathcal{G}_{B}\to\{0,1\}$ is graph-symmetric if $f(x,y)=f(x\circ\pi,y\circ\pi)$, where $x\in\mathcal{G}_{A},y\in\mathcal{G}_{B}$, and $\pi\in G_{n}$.
@@
A Facts 3.12 and 4.4 in Section 2
B Lemmas 3.7, 3.8 and 3.10 in Section 3
C Lemma 4.2 in Section 4
D Lemmas 5.1 and 5.2 in Section 5
@@ A.1 Pattern Matrix Method and Approximate Degree
For $f:\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$ and $10\leq\epsilon<$ 1, we say a real multilinear polynomial $p$ approximates $f$ with error $\epsilon$ if:
@@
$\Gamma(D)=\min\left\{|2k-n+1|:D(k),D(k+1)\neq*,D(k)\neq D(k+1)\text{ and }0\leq k\leq n-1\right\}.$
@@
$Q^{*}_{\delta}(F)$ $\geq\frac{1}{4}\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)\log\left(\frac{n}{t}\right)-\frac{1}{2}\log\left(\frac{3}{\epsilon-2\delta}\right).$
@@ A.2 Rank
We first restate and prove Fact 4.4 as follows: See 4.4
@@
By Example 2.12 in [32], $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)=\binom{n}{k}$. Since $2{\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}=J-\operatorname{\mathsf{DISJ}}$, where $J$ is the all-ones matrix, we have $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)\leq\operatorname{\mathrm{rank}}\left(J\right)+\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}\right)$. Thus, $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}\right)\geq\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)-\operatorname{\mathrm{rank}}\left(J\right)=\binom{n}{k}-1$. Since $\operatorname{\mathsf{EQ}}$ is an identity matrix, we have $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}\right)=\binom{n}{k}$. Similar to $\operatorname{\mathsf{DISJ}}$, we have $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{EQ}}\right)\geq\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}\right)-1=\binom{n}{k}-1$. ∎
@@ A.3 Approximate Rank
A.3.1 Error Reduction
@@
Let $10<E<$ 1 be given. Then for each interger $c\geq 1$, there exists a degree-$c$ real univariate polynomial $p(t)$ such that for any $1-E\leq|t|\leq 1+E$,
@@
Then we give A.7 and prove it as follow:
Fact A.7 (A generalized version of Corollary 1 in [30]).
Let $F$ be a matrix with $\left\{-1,1,*\right\}$ entries. Let $\epsilon,E$ be constants with $10<\epsilon<E<$ 1. Then $\log\operatorname{\widetilde{rank}}_{\epsilon}(F)=O\left(\log\operatorname{\widetilde{rank}}_{E}(F)\right)$.
@@
$28\sqrt{c}\left(1-\frac{(1-E)^{2}}{16}\right)^{c}\leq\epsilon/$ 2 .
@@
$\operatorname{\widetilde{rank}}_{\epsilon}(F)\leq\operatorname{\mathrm{rank}}(B)\leq(c+1)\operatorname{\mathrm{rank}}(A)^{c}=(c+1)\operatorname{\widetilde{rank}}_{E}(F)^{c}.$
@@
A.3.2 Lower Bound via Pattern Matrix Method
We first give some useful definitions, facts and lemmas.
@@ A.3.2 Lower Bound via Pattern Matrix Method
For any matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1, the $\epsilon$-approximate trace norm of $F$ is
$||F||_{\Sigma,\epsilon}\coloneqq\min_{A\in\mathcal{F}_{\epsilon}}||A||_{\Sigma},$
@@
For any incomplete matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1, we say a real matrix $A$ is a real representation of $F$ if:
@@
By Definitions A.10 and A.9, we have $||F||_{\Sigma,\epsilon}=\min_{A\in\mathcal{F}}||A||_{\Sigma,\epsilon}$.
@@
$||F||_{\Sigma,\epsilon}\geq\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\langle F,\Psi\rangle-\epsilon||\Psi||_{1}}{||\Psi||}.$
@@
Let $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1. Then
$||F||_{\Sigma,\epsilon}\geq\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\sum_{(x,y)\in\mathrm{dom}}\Psi_{x,y}F_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|-\epsilon||\Psi||_{1}}{||\Psi||}.$
@@
$||F||_{\Sigma,\epsilon}$ $=\min_{A\in\mathcal{F}}||A||_{\Sigma,\epsilon}$
$\geq\min_{A\in\mathcal{F}}\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\langle A,\Psi\rangle-\epsilon||\Psi||_{1}}{||\Psi||}$
$=\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\sum_{(x,y)\in\mathrm{dom}}\Psi_{x,y}F_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|-\epsilon||\Psi||_{1}}{||\Psi||}.$
@@
Let $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1. Then
$\operatorname{\widetilde{rank}}_{\epsilon}(F)\geq\frac{||F||_{\Sigma,\epsilon}^{2}}{\left(1+\epsilon\right)^{2}mn}.$
@@
$||F||_{\Sigma,\epsilon}=\min_{A\in\mathcal{F}_{\epsilon}}||A||_{\Sigma}\leq||A||_{F}\sqrt{\operatorname{\mathrm{rank}}{A}}\leq\sqrt{\sum_{i,j}A_{i,j}^{2}}\sqrt{\operatorname{\mathrm{rank}}{A}}\leq\sqrt{\left(1+\epsilon\right)^{2}mn}\sqrt{\operatorname{\mathrm{rank}}{A}}.$
@@
$\operatorname{\widetilde{rank}}_{\epsilon}(F)\geq\operatorname{\mathrm{rank}}{A}\geq\frac{||F||_{\Sigma,\epsilon}^{2}}{\left(1+\epsilon\right)^{2}mn}.$
@@
$||\Psi||_{1}$ $=1,$ (5)
$\sum_{(x,y)\in\mathrm{dom}}F_{x,y}\Psi_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|$ $>\epsilon,$
$||\Psi||$ $\leq\left(\frac{t}{n}\right)^{d/2}\left(2^{n+t}\left(\frac{n}{t}\right)^{t}\right)^{-1/2}.$
Finally, we give the following lower bound of approximate rank via pattern matrix method by A.15.
Fact A.15 (A generalized version of Theorem 1.4 in [40]).
@@
By Lemma A.12 and Equation 5, we have
$||F||_{\Sigma,\delta}\geq(\epsilon-\delta)\left(\frac{n}{t}\right)^{\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)/2}\sqrt{2^{n+t}(n/t)^{t}}.$ (6)
Since $F$ is a $2^{n}\times(n/t)^{t}2^{t}$ matrix, by Lemma A.13 and Equation 6, we have
$\operatorname{\widetilde{rank}}_{\delta}(F)$ $\geq\frac{||F||_{\Sigma,\delta}^{2}}{\left(1+\delta\right)^{2}2^{n+t}(n/t)^{t}}$
@@ A.4 Sampling in Communication Model
First, we give Fact A.16. Then we restate and prove its randomized version, Fact 3.12.
@@
The protocol of A.16 is a deterministic protocol based on binary search. Initially, the search space is $[n]$. In each round, Alice and Bob shrink the search space to one-half of the original search space. Finally, Alice and Bob find one index $i\in[n]$ such $x_{i}\neq y_{i}$ by $O(\log n)$ rounds. The deterministic protocol can be adjusted to a randomized protocol easily. By making the same random permutation to the bits of the inputs using public coins, Alice and Bob can find an index $i$ such that $x_{i}\neq y_{i}$ uniformly. ∎
@@ A.5 Quantum Amplitude Amplification
Given a unitary $U$ such that $U\ket{0}=\ket{\psi}$ and an orthogonal projector $\Pi$, there exists a quantum algorithm outputting an estimation $\tilde{p}$ of $p=\langle\psi|\Pi|\psi\rangle$ such that
@@
by $O\left(\sqrt{p}/\epsilon\right)$ calls to (the controlled versions of) $U,U^{{\dagger}}$ and $I-2\Pi$ with the bounded error.
As a direct corollary of A.17, we give the quantum version of 3.11 as follows.
@@
Given a set $S$, suppose there is a subset $S^{\prime}$ of $S$ such that $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$. Suppose we have a quantum sampler such that $\mathcal{S}\ket{0}=\frac{1}{\sqrt{|S|}}\sum_{i\in S}\ket{i}$. Let orthogonal projector $\Pi$ be defined as $\Pi=\sum_{i\in S^{\prime}}\ket{i}\bra{i}$ |. Since $\frac{|S^{\prime}|}{|S|}=\bra{0}\mathcal{S}^{{\dagger}}\Pi\mathcal{S}\ket{0}$, we can decide whether $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$ by $O\left(\sqrt{\beta}/\epsilon\right)$ calls of (the controlled versions of) $\mathcal{S},\mathcal{S}^{{\dagger}}$ and $I-2\Pi$ with the bounded error.
@@ B.1 The Lower Bound on Quantum Communication Complexity
In this section, we restate and prove the following lemmas: See 3.7
@@
    Suppose there exists a quantum protocol $\mathcal{P}$ that solves $\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$. To solve $\mathsf{ESetInc}_{a,b,c,g}^{n}$, Alice and Bob can append some bit strings to their input, respectively. First, Alice appends $\ell_{1}$ 1’s to her input string, and Bob appends $\ell_{1}$ 0’s in the same positions. Second, Bob appends $\ell_{2}$ 1’s to his input, and Alice appends $\ell_{2}$ 0’s in the same positions. Third, Alice and Bob append $\ell_{3}$ 1’s at the same time. Finally, Alice and Bob run the protocol for $\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$.
@@
    Suppose there exists a quantum protocol solving $\mathsf{ESetInc}_{a,b,c,g}^{n}$. To solve $\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}$, Bob can first flip each bit of his input, and then Alice and Bob run the protocol for $\mathsf{ESetInc}_{a,b,c,g}^{n}$. Thus $Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)\leq Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)$. By a similar argument, we have
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)$ $\leq Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right),$
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)$ $\leq Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right),$
    $Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$ $\leq Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right).$
    Thus, $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$.
@@
    Suppose there exists a quantum protocol solving $\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}$. To solve $\mathsf{ESetInc}_{a,b,c,g}^{n}$, Alice and Bob can repeat their input strings $k$ times and run the protocol for $\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}$.
See 3.8
@@
Let $P$ be the $\left(2k,k,f_{k,l}\right)$-pattern matrix defined as Definition A.1. A.4 implies that $Q^{*}_{1/3}\left(P\right)=\Omega\left(\sqrt{kl}\right)$, where
@@
For any $x\in\{0,1\}^{2k}$, we have $x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}\in\{0,1\}^{4k}$ and $|x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}|=2k$; for any $V\in\mathcal{V}(4k,k)$, we have $V\in\{0,1\}^{4k}$ and $|V|=k$. Thus, $P$ is a submatrix of $\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}$, which is defined as
$\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}(x,y)\coloneqq\begin{cases}-1,&\text{if $|x|=2k,|y|=k$ and $|x\land y|=l-1/2$},\\ 1,&\text{if $|x|=2k,|y|=k$ and $|x\land y|=l+1/2$},\\ *,&\text{otherwise.}\end{cases}$
As a result, we have $Q^{*}_{1/3}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\geq Q^{*}_{1/3}\left(P\right)=\Omega\left(\sqrt{kl}\right)$. ∎

### v2

@@ 6 Conclusion
This paper proves that the randomized communication complexity of permutation-invariant Boolean functions is at most quadratic of the quantum communication complexity (up to a polylogarithmic factor of the input size). Our results suggest that symmetries prevent exponential quantum speedups in communication complexity, extending the analogous research on query complexity. Furthermore, we prove that the Log-rank Conjecture and Log-approximate-rank Conjecture hold for non-trivial permutation-invariant Boolean functions (up to a polylogarithmic factor of the input size). There are some interesting problems to explore in the future.
  • Permutation invariance over higher alphabets. In this paper, the permutation-invariant function is a binary function. The interesting question is to generalize our results to larger alphabets, i.e., to permutation-invariant functions of the form $f:\left\{0,1,...,m\right\}^{n}\times\left\{0,1,...,m\right\}^{n}\to\left\{-1,1,*\right\}$ where $m\in\mathbb{N}$ and $m>1$.
  • Generalized permutation invariance. It is possible to generalize our results for a larger class of symmetric functions. One candidate might be a class of functions that have graph-symmetric properties. Suppose $\mathcal{G}_{A},\mathcal{G}_{B}$ are two sets of $n$-vertices graphs, and $G_{n}$ is a group that acts on the edges of an $n$-vertices graph and permutes them in a way that corresponds to relabeling the vertices of the underlying graph. A function $f:\mathcal{G}_{A}\times\mathcal{G}_{B}\to\{-1,1,*\}$ is graph-symmetric if $f(x,y)=f(x\circ\pi,y\circ\pi)$, where $x\in\mathcal{G}_{A},y\in\mathcal{G}_{B}$, and $\pi\in G_{n}$.
@@
A Facts 3.11 and 4.3 in Section 2
B Lemmas 3.6, 3.7 and 3.9 in Section 3
C Lemmas 5.1 and 5.2 in Section 5
D Communication complexity of Gap-Hamming-Distance
@@ A.1 Pattern Matrix Method and Approximate Degree
For $f:\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$ and $0\leq\epsilon<1$, we say a real multilinear polynomial $p$ approximates $f$ with error $\epsilon$ if:
@@
$\Gamma(D)$ $=\min\{|2k-n+1|:D(k),D(k+1)\neq*,$
$D(k)\neq D(k+1)\text{ and }0\leq k\leq n-1\}.$
@@
$Q_{\delta}(F)$ $\geq\frac{1}{4}\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)\log\left(\frac{n}{t}\right)-\frac{1}{2}\log\left(\frac{3}{\epsilon-2\delta}\right).$
@@ A.2 Rank
We first restate and prove Fact 4.3 as follows: See 4.3
@@
By Example 2.12 in [31], $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)=\binom{n}{k}$. Since $2{\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}=J-\operatorname{\mathsf{DISJ}}$, where $J$ is the all-ones matrix, we have $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)\leq\operatorname{\mathrm{rank}}\left(J\right)+\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}\right)$. Thus, $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}\right)\geq\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)-\operatorname{\mathrm{rank}}\left(J\right)=\binom{n}{k}-1$. Since $\operatorname{\mathsf{EQ}}^{\prime}$ is an identity matrix, we have $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}\right)=\binom{n}{k}$. Similar to $\operatorname{\mathsf{DISJ}}$, we have $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{EQ}}\right)\geq\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}\right)-1=\binom{n}{k}-1$. ∎
@@ A.3 Approximate Rank
First, utilizing Facts A.5 and A.6, we prove A.7 to show the error reduction of approximate rank.
@@
Let $0<E<1$ be given. Then for each interger $c\geq 1$, there exists a degree-$c$ real univariate polynomial $p(t)$ such that for any $1-E\leq|t|\leq 1+E$,
@@
Fact A.7 (Error reduction, a generalized version of Corollary 1 in [29]).
Let $F$ be a matrix with $\left\{-1,1,*\right\}$ entries. Let $\epsilon,E$ be constants with $0<\epsilon<E<1$. Then $\log\operatorname{\widetilde{rank}}_{\epsilon}(F)=O\left(\log\operatorname{\widetilde{rank}}_{E}(F)\right)$.
@@
$8\sqrt{c}\left(1-\frac{(1-E)^{2}}{16}\right)^{c}\leq\epsilon/2.$
@@
$\operatorname{\widetilde{rank}}_{\epsilon}(F)$ $\leq\operatorname{\mathrm{rank}}(B)$
$\leq(c+1)\operatorname{\mathrm{rank}}(A)^{c}$
$=(c+1)\operatorname{\widetilde{rank}}_{E}(F)^{c}.$
@@
Next, we first give some useful definitions, facts, and lemmas about the pattern matrix method [40].
@@
For any matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $0\leq\epsilon<1$, the $\epsilon$-approximate trace norm of $F$ is
[…]
@@
For any incomplete matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $0\leq\epsilon<1$, we say a real matrix $A$ is a real representation of $F$ if:
@@
[…]
@@
[…]
@@
Let $F\in\left\{-1,1,*\right\}^{m\times n}$ and $0\leq\epsilon<1$. Then for any $\Psi\in\mathbb{R}^{m\times n}$ such that $\Psi||\neq 0$, we have
[…]
is a lower bound of $F||_{\Sigma,\epsilon}$, where $\mathrm{dom}:=\{(x,y):F_{x,y}\in\{-1,1\}\}$.
@@
[…]
@@
Let $F\in\left\{-1,1,*\right\}^{m\times n}$ and $0\leq\epsilon<1$. Then
[…]
@@
[…]
$\leq\sqrt{\sum_{i,j}A_{i,j}^{2}}\sqrt{\operatorname{\mathrm{rank}}{A}}$
$\leq\sqrt{\left(1+\epsilon\right)^{2}mn}\sqrt{\operatorname{\mathrm{rank}}{A}}.$
@@
[…]
@@
[…]
$\epsilon$ $<\sum_{(x,y)\in\mathrm{dom}}F_{x,y}\Psi_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|,$
[…]
where $\mathrm{dom}:=\{(x,y):F_{x,y}\in\{-1,1\}\}$.
Finally, we give the following lower bound of approximate rank via the pattern matrix method by Lemma A.15.
Lemma A.15 (A generalized version of Theorem 1.4 in [40]).
@@
By Lemma A.12 and (8), we have
$F||_{\Sigma,\delta}\geq(\epsilon-\delta)\left(\frac{n}{t}\right)^{\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)/2}\sqrt{2^{n+t}(n/t)^{t}}.$ (9)
Since $F$ is a $2^{n}\times(n/t)^{t}2^{t}$ matrix, by Lemma A.13 and (9), we have
[…]
@@ A.4 Sampling in Communication Model
First, we give Fact A.16. Then we restate and prove its randomized version, Fact 3.11.
@@
The protocol of A.16 is a deterministic protocol based on binary search. Initially, the search space is $\left\{0,1,...,n-1\right\}$. In each round, Alice and Bob shrink the search space to one-half of the original search space. Finally, Alice and Bob find one index $i\in\left\{0,1,...,n-1\right\}$ such $x_{i}\neq y_{i}$ by $O(\log n)$ rounds. The deterministic protocol can be adjusted to a randomized protocol easily. By making the same random permutation to the bits of the inputs using public coins, Alice and Bob can find an index $i$ such that $x_{i}\neq y_{i}$ uniformly. ∎
@@ A.5 Quantum Amplitude Amplification
[…]
@@
by $O\left(\sqrt{p}/\epsilon\right)$ calls to (the controlled versions of) $U,U^{{\dagger}}$ and $I-2\Pi$ with bounded error.
As a direct corollary of A.17, we give the quantum version of 3.10 as follows.
@@
[…]
@@ B.1 The Lower Bound on Quantum Communication Complexity
In this section, we restate and prove the following lemmas: See 3.6
@@
    Suppose there exists a quantum protocol $\mathcal{P}$ that solves $\mathrm{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$. To solve $\mathrm{ESetInc}_{a,b,c,g}^{n}$, Alice and Bob can append some bit strings to their input, respectively. First, Alice appends $\ell_{1}$ 1’s to her input string, and Bob appends $\ell_{1}$ 0’s in the same positions. Second, Bob appends $\ell_{2}$ 1’s to his input, and Alice appends $\ell_{2}$ 0’s in the same positions. Third, Alice and Bob append $\ell_{3}$ 1’s at the same time. Finally, Alice and Bob run the protocol for $\mathrm{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$.
@@
    Suppose there exists a quantum protocol solving $\mathrm{ESetInc}_{a,b,c,g}^{n}$. To solve $\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}$, Bob can first flip each bit of his input, and then Alice and Bob run the protocol for $\mathrm{ESetInc}_{a,b,c,g}^{n}$. Thus $Q\left(\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}\right)\leq Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)$. By a similar argument, we have
    $Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)$ $\leq Q\left(\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}\right),$
    $Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)$ $\leq Q\left(\mathrm{ESetInc}_{n-a,b,b-c,g}^{n}\right),$
    $Q\left(\mathrm{ESetInc}_{n-a,b,b-c,g}^{n}\right)$ $\leq Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right).$
    Thus, $Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)=Q\left(\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}\right)=Q\left(\mathrm{ESetInc}_{n-a,b,b-c,g}^{n}\right)$.
@@
    Suppose there exists a quantum protocol solving $\mathrm{ESetInc}_{ka,kb,kc,kg}^{kn}$. To solve $\mathrm{ESetInc}_{a,b,c,g}^{n}$, Alice and Bob can repeat their input strings $k$ times and run the protocol for $\mathrm{ESetInc}_{ka,kb,kc,kg}^{kn}$.
See 3.7
@@
Let $P$ be the $\left(2k,k,f_{k,l}\right)$-pattern matrix defined as Definition A.1. A.4 implies that $Q_{1/3}\left(P\right)=\Omega\left(\sqrt{kl}\right)$, where
@@
For any $x\in\{0,1\}^{2k}$, we have $x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}\in\{0,1\}^{4k}$ and $|x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}|=2k$; for any $V\in\mathcal{V}(4k,k)$, we have $V\in\{0,1\}^{4k}$ and $|V|=k$. Thus, $P$ is a submatrix of $\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}$, which is defined as
$\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}(x,y)$
$\coloneqq\begin{cases}-1,&\text{if $|x|=2k,|y|=k$ and $|x\land y|=l-1/2$},\\ 1,&\text{if $|x|=2k,|y|=k$ and $|x\land y|=l+1/2$},\\ *,&\text{otherwise.}\end{cases}$
As a result, we have $Q_{1/3}\left(\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}\right)\geq Q_{1/3}\left(P\right)=\Omega\left(\sqrt{kl}\right)$. ∎